OpenStudy (anonymous):

Which shows the expression below simplified? (9 × 104) × (9 × 10-4) 8.1 × 100 8.1 × 101 8.1 × 10-1 8.1 × 102

OpenStudy (anonymous):

The numbers after the 10s are the exponents.

OpenStudy (anonymous):

K ! we have : 9 * 10000 =90000 9*10^-4=9*0.0001=0.0009 Then : can say : 9*9=81 or 8.1*10 Got it ?

OpenStudy (anonymous):

Yes, thank you! :D

OpenStudy (anonymous):

Er, how do I find the exponent at the end of the answer? o:

OpenStudy (anonymous):

The answer is B .!

OpenStudy (anonymous):

Oh, okay. Thanks!

OpenStudy (anonymous):

OH! Because the exponent is only one! I got you. Thanks! :D

OpenStudy (anonymous):

;)